OpenStudy (anonymous):

lexi has planted seeds in 3/5 of the garden. she used 1/2 pound of seeds. how many pounds will she use for the entire garden

OpenStudy (anonymous):

1/4 ?

OpenStudy (jdoe0001):

so if 3/5 is just a fraction of the garden, 5/5 will be the entire garden itself, so \(\large \begin{array}{ccll} seeds&garden\\ \hline\\ \frac{1}{2}&\frac{3}{5}\\ \quad \\ x&\frac{5}{5} \end{array}\implies \cfrac{\frac{1}{2}}{x}=\cfrac{\frac{3}{5}}{\frac{5}{5}}\)

OpenStudy (anonymous):

so would the answer be \[\frac{ 1 }{ 2 }\] ? @jdoe0001

OpenStudy (jdoe0001):

hmm well... can't be, she already used 1/2 of the seeds for 3/5 of the garden so it has to be more, since she'd need more seeds for the other 2/5 of the garden so it'd be 1/2 + SOME AMOUNT

OpenStudy (anonymous):

for some reason i got \[1 \frac{ 1 }{5 }\]

OpenStudy (jdoe0001):

let's see gimme one sec

OpenStudy (anonymous):

this is how i set up my equation: \[\frac{ 3 }{ 5 }\div \frac{ 1 }{ 2 } =\frac{ 6 }{5 } which simplifies \to 1\frac{ 1 }{ 5}\]

OpenStudy (jdoe0001):

..... yeap \( \large{\begin{array}{ccll} seeds(lbs)&garden\\ \hline\\ \frac{1}{2}&\frac{3}{5}\\ \quad \\ x&\frac{5}{5} \end{array}\implies \cfrac{\frac{1}{2}}{x}=\cfrac{\frac{3}{5}}{\frac{5}{5}} \\ \quad \\ \textit{cross-multiplying we get }x=\cfrac{\frac{5}{5}\cdot \frac{1}{2}}{\frac{3}{5}}\implies x=\cfrac{1\cdot \frac{1}{2}}{\frac{3}{5}} \\ \quad \\ x=\cfrac{\frac{1}{2}}{\frac{3}{5}}\implies x=\cfrac{1}{2}\cdot \cfrac{5}{3}}\)

OpenStudy (anonymous):

so the answer would be \[\frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

sorry but website is a bit lagged yeap 5/6